OpenStudy (anonymous):

PLEASE HELP, For a rectangular prism, if the base length was shrunk by a factor of two-thirds, the width was quadrupled, and height was cut in half, how would the changes affect the surface area formula?

OpenStudy (anonymous):

Heres the problem again with answers.

OpenStudy (mathstudent55):

The total surface area of a rectangular prism is: SA = 2LW + 2LH + 2WH Now take the length, L, and replace it with 2L/3. Take the width, W, and replace it with 4W. Take the height, H, and replace it wirth H/2. Make all thses substitutions in the formula above and figure out the result.

OpenStudy (anonymous):

SA=2L/3+2L2+24/2 LIKE THAT?

OpenStudy (anonymous):

Im not getting your instructions..

OpenStudy (mathstudent55):

L is substituted by 2L/3. The first term of the surface area is 2LW. L now becomes 2L/3, and W now becomes 4W, so 2LW becomes 2 * (2L/3) * (4W)

OpenStudy (anonymous):

im still confused..

OpenStudy (anonymous):

So my base is 1/16 ?

OpenStudy (mathstudent55):

\(SA = 2~~~~~L~~~~~~W ~~~~+2~~~~~L~~~~~~~~H~~+~2~~~W~~~~~~H\) \(SA = 2(2L/3)(4W) + 2(2L/3)(H/2) + 2(4W)(H/2)\)

OpenStudy (anonymous):

So is D my answer?

OpenStudy (mathstudent55):

\(SA = \frac{16}{3} LW + \frac{2}{3}LH + 4WH\)

OpenStudy (anonymous):

C. Thanks.

OpenStudy (mathstudent55):

That's it. Wlcm.